Please explain for each stepA company borrows $17,000 to purchase a new piece of equipment. The loan will be repaid in one lump sum at the end of 5 years. The bank offers to loan the money at 0.5% per month, but the company prefers to repay the loan at 6% per year. If the company is successful at getting the bank to agree to its preferred terms, how much will the company save in interest on the loan? Express your answer in $ to the nearest whole $.
Answer #1

So, here “$17,000” be the principal with “6%” interest rate compounded annually, => after “5 year” the total amount will be.

=> P*(1+r)^5 = 17,000*1.06^5 = 22,749.8348 = $22,750.

Now, if the interest rate is “0.5%” compounded monthly then the amount will be.

=> $17,000*1.005^(5*12)= $17,000*1.005^60 = 22,930.4526 = $22,930.

So, if the company is successful at getting the bank to agree to its preferred terms, => the company will save, “$22,930 – $22,750 = $180”.